The events A and B are mutually exclusive. If P(A) = 0.4 and P(B) = 0.3, what is P(A or B)?

Answer by ikleyn(52878)

.

If events are mutually exclusive, then  P( A or B) = P(A) + P(B) = 0.4 + 0.3 = 0.7.    ANSWER

Solved.